myininaya (myininaya):

no one answer it only satellite can solve for a+b a^2+2ab+b^2+6a+6b+9=0

myininaya (myininaya):

a^2+2ab+b^2+6a+6b+9=0 for prettiness :)

myininaya (myininaya):

oops didnt use latex

OpenStudy (anonymous):

Can I ask, out of curiosity, why only satellite can answer?

myininaya (myininaya):

\[a^2+2ab+b^2+6a+6b+9=0\]

myininaya (myininaya):

because its his challenge question do you want one?

myininaya (myininaya):

lol

OpenStudy (anonymous):

I kinda want this one, since I know how to do it...but I'll take another one, sure.

OpenStudy (anonymous):

\[((a+b)+3)^2\]?

myininaya (myininaya):

ok blue you can steel it i guess

myininaya (myininaya):

go blue!

OpenStudy (anonymous):

\[(a+b)^{2}+6(a+b)+9=0\] Let x=a+b We have \[x ^{2}+6x+9=0\] Leading to x=-3....so a+b=-3

myininaya (myininaya):

very good blue

OpenStudy (anonymous):

I owe satellite an apology, I suppose?

myininaya (myininaya):

lol

OpenStudy (anonymous):

You can give him my question :)

myininaya (myininaya):

the problem was to easy i should give him a harder one

myininaya (myininaya):

help me think blue

OpenStudy (anonymous):

Well, that makes me feel a lot better!

OpenStudy (anonymous):

just got here. am i supposed to solve for a? b?

myininaya (myininaya):

a+b

myininaya (myininaya):

i seen this question on the gre and i was like omg what and then i was like oh its just a quadratic

OpenStudy (anonymous):

omg this is like factoring sort of like the one is sent you yes?

OpenStudy (anonymous):

it is a perfect square?

myininaya (myininaya):

(a+b)^2+6(a+b)+9=0 yes perfect square (a+b+3)^2=0 a+b+3=0 a+b=-3

myininaya (myininaya):

blue i meant to say its a tough one lol do you feel better now?

OpenStudy (anonymous):

wow nice . \[\color{blue}{\text{perfect square}}\]

OpenStudy (anonymous):

Yes, much :)

myininaya (myininaya):

:)

OpenStudy (anonymous):

\[\color{blue}{\text{go blue!}}\]

myininaya (myininaya):

that kindof looks purple

OpenStudy (anonymous):

\[\color{pink}{\text{my my my myininaya}}\]

OpenStudy (anonymous):

http://www.youtube.com/watch?v=kVdnqEyToqg

myininaya (myininaya):

my sharona!

myininaya (myininaya):

love that song

myininaya (myininaya):

oops male i didn't mean to leave you out i didnt notice what you wrote good job

OpenStudy (anonymous):

perfect blend of pre - new wave splashy drums anthem rock and proto new wave spare riffs